OpenStudy (nicole143):

Simplify the complex fraction. 2-[2/x] / 3-[1/x] I thought I had this but came down to the wrong answer. Can you help me? @mathmale

OpenStudy (mathmale):

Hello, N! Just wanted to be certain I undrstand what you're asking. Do you agree with the following interpretation?\[\frac{ 2-(\frac{ 2 }{ x }) }{3-(\frac{ 1 }{ x }) }?\]

OpenStudy (nicole143):

Yes(:

OpenStudy (nicole143):

Do you want to see what I have? Maybe I don't have to start all the way over ha

OpenStudy (mathmale):

Why sure! If you have a quick way to upload an image of your work, please do that.

OpenStudy (nicole143):

Sorry for the writing ha

OpenStudy (isaiah.feynman):

I don't know this method you used that involves adding y. What I'd do is just add the fraction and the whole number up, then the fraction and the whole number down. Finally dividing.

OpenStudy (nicole143):

I would have to see what you are saying because I don't get it..

OpenStudy (nicole143):

The adding of "xy" was me trying to follow my book

OpenStudy (mathmale):

Nicole, my suggestion is that you multiply both numerator and denom. of the given rational fraction by ONLY x. We can deal with that "xy" later if need be. Please, would you go ahead and mult the num. by x now and type in the result (here)?

OpenStudy (nicole143):

Yes, one second.

OpenStudy (nicole143):

2x - 2 ?

OpenStudy (mathmale):

Or 2(x-1). Perfect. Now mult the den. by x and type the res. here.

OpenStudy (nicole143):

3(x-1) ?

OpenStudy (mathmale):

Here's what I see: x[3-1/x) = 3x -1 (not 3[x-1]. What do you think?

OpenStudy (nicole143):

Can you do it in the equation thing so I can see it?

OpenStudy (nicole143):

Just the first one..

OpenStudy (mathmale):

I'm going to take your numerator and denominator and write them in Equation Editor ('tho I'm not positive this is what you want to see).\[\frac{ 2(x-1) }{ x(3-[\frac{ 1 }{ x }] )}\]

OpenStudy (mathmale):

If you complete the mult. in the den., you'll get 3x -1. Are you comfortable with that?

OpenStudy (nicole143):

Yes, I see it now

OpenStudy (mathmale):

That's great! I need to get off my computer now, but will be available on and off over the next few days. Great to "see" you! MM

OpenStudy (nicole143):

Oh, Okay. Thank you very much as always! Have a good night. -Nicole

OpenStudy (mathmale):

Bye, Nicole!!